# Revisiting the Jaynes–Cummings model with time-dependent coupling

### v1

@@ Abstract
The Jaynes–Cummings (JC) model stands as a fully quantized, fundamental framework for exploring light–matter interactions, a timely reflection on a century of quantum theory. The time-dependent Jaynes–Cummings (TDJC) model introduces temporal variations in certain parameters, which often require numerical methods. However, under the resonance condition, exact solutions can be obtained, offering insight into a variety of physical scenarios. In this work, we study the resonant TDJC model considering different modulations of the atom–field coupling. The model is presented and an analytical solution derived in a didactic way, allowing us to examine how time-dependent couplings affect atomic population inversion and atom–field entanglement. We also consider an atom traversing a partially cooled cavity, which induces periodicity and reveals the combined effects of atomic motion and thermal fluctuations. The Bloch vector is used to analyze the dynamics of the system, including the atomic state purity, and reveals phenomena such as atomic dipole alignment with the field due to the oscillating coupling, as well as atomic population trapping, which arises by increasing the initial mean thermal photon number.
@@ I Introduction
In 2025, we mark the centenary of the inception of quantum mechanics, a revolutionary theory that continues to drive progress through the ongoing development of quantum technologies. The birth of quantum theory is deeply connected to the investigation of electromagnetic (EM) radiation and atoms, which remain central to both foundational investigations and emerging applications. More recently, the development of the laser, a coherent light source, gave rise to the subfield of nonrelativistic quantum electrodynamics theory known as quantum optics, which basically explores the nature and effects of quantized light. In this context, the most elementary quantum model of light–matter interaction consists of a highly idealized atom (modeled as a two-level system) coupled to a single mode of the quantized EM field. Under the rotating wave approximation (RWA), this model, known as the Jaynes–Cummings (JC) model [1, 2], remarkably admits an exact analytical solution, and despite the simplifying assumptions underlying its construction [3], the JC model provides deep insights into several fundamental aspects of light–matter interaction. Originally developed to compare semiclassical and quantum approaches to spontaneous emission [4], the JC model has since evolved to uncover nonclassical phenomena such as the experimentally observed collapses [5] and revivals of the Rabi oscillations (RO) [6, 7, 8, 9, 10], squeezing [11, 12], sub-Poissonian statistics [13, 14], atom-field entanglement [15, 16, 17], and the generation of Schrödinger cat states [18, 19]. Furthermore, the core of the model is a two-level system (TLS) coupled to a harmonic oscillator, which allows it to be mapped onto trapped ion [20, 21, 22] and relativistic systems [23]. Regarding applications, the JC framework is employed in several quantum information processing schemes [24, 25]. There are also generalizations of the model involving additional field modes and atomic levels [26, 27], multiple two-level atoms [28, 29], quantum deformations [30, 31, 32], Kerr-like media [33, 34] and external pumpings [35, 36] Besides, the model naturally extends to the description of open quantum systems, where interaction with the environment leads to dissipation and decoherence [37, 38, 39].
@@
A particularly interesting generalization is obtained when the model parameters are allowed to vary in time, giving rise to the time-dependent JC (TDJC) model [40, 24]. This may account for atomic motion [41, 42], transient effects in the cavity [43, 44], or the incorporation of stochastic aspects [45]. However, unlike the standard JC model, analytical methods are not always applicable. Notable exceptions are the Nikitin [43], Rosen-Zener [44], Landau-Zener [46, 47] and Demkov–Kunike [46] models. Remarkably, under the resonance condition, it is possible to obtain exact solutions considering a time-dependent atom-field coupling [41, 48, 42]. From an experimental point of view, Ref. [49] suggests that a time-dependent coupling may be achievable in cavity quantum electrodynamics through variations in the atom’s position or by changing an external electric field.
@@
The objective of this work is twofold: we first present a didactic overview of the solution of the resonant TDJC model in terms of the evolution operator. We discuss two specific cases of time-dependent atom–field coupling: (i) linear [48] and (ii) hyperbolic secant [44]. The latter is particularly desirable in quantum control [50], where manipulating the TLS is one potential objective [51, 52]. We focus on the dynamics of the atomic population inversion and the von Neumann entropy, examining the impact of the time dependence on both quantities. The subsystems are assumed to be initially prepared in pure states, in which case the entropy directly reflects the atom–field entanglement. In the second part of the paper, we consider the atom moving inside the cavity, which leads to a sinusoidal time dependence for the atom–field coupling. Related scenarios have been explored in the literature, particularly in studies of entanglement [53, 54], self-induced transparency effects [41], and the role of atomic coherence in the JC model with motion [55]. We also assume a different set of initial conditions, viz., the field in a thermal state and the atom prepared either in the excited state or in a superposition of atomic ground and excited states (eigenstate of $\hat{\sigma}_{x}$). The JC model with a partially cooled cavity is relevant, for example, to the study of entanglement [56, 54], and is also experimentally accessible [9]. We analyze the atomic evolution from a different perspective, focusing on the Bloch vector dynamics and compare our results to the established findings [57, 56, 58, 59] within the context of constant coupling. We find that the behavior of the studied quantities, basically the components and modulus of the Bloch vector, is generally dominated by the periodic coupling, even in the presence of thermal noise, in contrast to the constant coupling case [56]. We note that a periodic evolution of both the field entropy and the atomic inversion, induced by atomic motion and in the absence of thermal noise, was reported in [42]. Other effects may arise, manifesting in the Bloch vector dynamics. Two control parameters in the system, $p$, the number of half-wavelengths of the standing-wave cavity, and $\zeta_{3}$, proportional to the velocity of the moving atom, have similar effects. For instance, smaller values of $p$ and $\zeta_{3}$ lead to a dynamical evolution with longer periods, more closely resembling the behavior of the constant-coupling case. Interestingly, increasing the thermal fluctuations, which corresponds to a larger mean photon number in the initial field state, leads to a form of atomic population trapping [57], while the periodicity of the Bloch vector dynamics tends to persist despite the added noise.
@@ III Time-dependent coupling dynamics
In this work, we are interested in how time-dependent coupling parameters affect the on-resonance dynamics of the JC model. Other works that adopt a similar approach can be found in Refs. [41, 90, 48, 42, 91].
@@ III.1 Linear
Joshi and Lawande [48] considered the linear coupling modulation, given by
@@
where $\lambda_{0}$ and $\zeta_{1}$ are control parameters that determine whether the change in the coupling is either sudden or adiabatic. According to Ref. [24], this modulation can be used to represent the scenario of a well-localized atom in a spatially varying mode. The authors observed that, for an initial coherent field state, this time-dependent coupling alters the revival time. More recently, Hu and Tan [92] studied a double JC model with linear modulation, focusing on the disappearance of entanglement between the two atoms. They suggested that a linear coupling parameter could be realized in cavity quantum electrodynamics by controlling the intensity of an external electric field.
@@ III.2 Hyperbolic Secant
This type of coupling, can be used to simulate transient effects in the cavity, allowing the interaction to be switched on and off depending on the choice of the initial time. The population inversion in this scenario is given by
@@
In Fig. 5, we observe how this coupling, which decays to zero, can lead to different asymptotic values for both population inversion and entanglement. In Fig. 5(a), we set $\zeta_{2}=0.3$ (faster decay of coupling), causing the dynamics to halt before reaching the revival time. In contrast, in Fig. 5(b), we adopt a slower decay for the coupling. In both cases, the dynamics cease after the collapse, causing the asymptotic value of $W(t)$ to be 0. However, for another combination of parameters, it would be possible to obtain different asymptotic values. Furthermore, another important aspect is that, even though the asymptotic value of the population inversion is the same in both cases, the von Neumann entropy value is significantly different. The coherence terms of the density matrix, which do not affect the atomic probabilities, are influenced by the different atom-field decoupling times, and this is reflected in the entanglement behavior.
@@ IV Atomic motion
As mentioned in the previous section, Schlicher was the first to introduce a time-dependent coupling parameter. In that context, he employed a sinusoidal modulation to account for the motion of the atom along the standing-wave [41]. Since then, this type of time dependence has been extensively studied: Bartzis investigated squeezing in the cavity degree of freedom [90], Wilkens and Meystre extended the approach to maser systems [93], and Fang computed the von Neumann entropy and the Husimi function [42], considering an initial coherent state of the cavity. Other works regarding atomic motion in the model can be found in Refs. [94, 95, 96, 97, 98]. In this case, we adopt a semiclassical approach, in which the atomic motion is treated classically—an approximation justified when the atom’s kinetic energy significantly exceeds the interaction energy of the JC model [24].
@@ IV.1 Bloch vector
Figure 8: Trajectories and time-evolution of the Bloch vector. In panels (a) and (e), atomic motion is not considered, and we set the parameter $\lambda=1$. Conversely, in the remaining panels, atomic motion is included, with parameters $\lambda_{0}=1$, $v=\zeta_{3}\pi/L$ and considered a time interval $0<t<40$. In panels (b), (c), (f) and (g), we set $p=1$, whereas in panels (d) and (h), we use $p=2$. In (b) and (f), we employ $\zeta_{3}=0.25$, while in the other panels, we fix $\zeta_{3}=1$. In all cases, the initial average photon number is $\langle n\rangle=0.5$. In panels (a), (b), (c) and (d), we present a parameter plot of the $R_{x}(t)$ and $R_{z}(t)$ components as a function of time. We observe that atomic motion dominates the disordered behavior of the Bloch vector. Panels (e), (f), (g) and (h) depict the time evolution of $R_{x}(t)$ (solid green line), $R_{z}(t)$ (dotted purple line), and $R(t)$ (dashed black line). The periodicity once again reveals the presence of sinusoidal coupling. Additionally, an increase in $p$ diminishes the amplitude of $R_{x}(t)$ and $R(t)$.
@@
In Fig. 8, we present the motion of the Bloch vector for both constant and sinusoidal coupling scenarios. In the constant coupling case, shown in Fig. 8(a) and (e), we set $\lambda=1$. For the sinusoidal coupling case, depicted in the other panels of Fig. 8, we fix the parameters as $\lambda_{0}=1$ and $v=\zeta_{3}L/\pi$. In panels (b), (c), (f) and (g), we set $p=1$, while in panels (d), and (h), we set $p=2$. In (b) and (g), we employ $\zeta_{3}=0.25$, while in the other panels, we fix $\zeta_{3}=1$. At the outset, we choose a low average photon number, $\langle n\rangle=0.5$, corresponding to a low temperature, for all scenarios. Comparing panels (a) and with (b), (c) and (d), we observe that the sinusoidal coupling introduces periodicity into the Bloch vector’s motion, replacing the disordered behavior. In contrast, for a smaller $\zeta_{3}$, as shown in (b), the behavior partially resembles that of the time-independent coupling case, exhibiting reduced-amplitude oscillations within each period. This suggests that for lower values of the parameter $\zeta_{3}$, an intermediate regime emerges in which vestiges of the aperiodic evolution persist. In Figs. 8(f), (g) and (h), which show the components and absolute value of the Bloch vector, we again observe the periodicity inherited from the modulated coupling, as compared to Fig. 8(e). Furthermore, as shown in the panels where atomic motion is considered, an increase in $p$ and $\zeta_{3}$ results in a higher frequency of the oscillations, while the amplitudes of the oscillations in $R_{x}(t)$ and $R(t)$ decrease. As a result, the atom tends to be more aligned with the field, leading to a higher time-averaged purity of the atomic state, whereas $R_{z}(t)$ oscillates with approximately the same amplitude. In all examples with atomic motion, we begin with $R(0)=R_{x}(0)=1$ and $R_{z}(0)=0$, meaning that initially, the atomic subsystem is pure, the atomic transition is aligned with the field and the excited and ground populations are equal. This behavior is a direct consequence of the trigonometric dependence of the coupling area $A(t)$, as discussed previously, which can also be interpreted in terms of Bloch vector rotations for the chosen velocity, as predicted by Joshi and Xiao [55].

### v2

@@ Abstract
The Jaynes–Cummings (JC) model stands as a fully quantized, fundamental framework for exploring light–matter interactions, a timely reflection on a century of quantum theory. The time-dependent Jaynes–Cummings (TDJC) model introduces temporal variations in certain parameters, which often require the use of numerical methods. However, under the resonance condition, exact solutions can be obtained, offering insight into a variety of physical scenarios. In this work, we study the resonant TDJC model considering different modulations of the atom–field coupling. The model is presented and an analytical solution derived in a didactic way, allowing us to examine how time-dependent couplings affect atomic population inversion and atom–field entanglement. We also consider an atom traversing a partially cooled cavity, which induces periodicity and reveals the combined effects of atomic motion and thermal fluctuations. The Bloch vector is used to analyze the dynamics of the system, including the atomic state purity, and reveals phenomena such as atomic dipole alignment with the field due to the oscillating coupling, as well as atomic population trapping, which arises by increasing the initial mean thermal photon number.
doi: 10.1007/s13538-025-01949-w
@@ I Introduction
In 2025, we mark the centenary of the inception of quantum mechanics, a revolutionary theory that continues to drive progress through the ongoing development of quantum technologies. The birth of quantum theory is deeply connected to the investigation of electromagnetic (EM) radiation and atoms, which remain central to both foundational investigations and emerging applications. More recently, the development of the laser, a coherent light source, gave rise to the subfield of nonrelativistic quantum electrodynamics theory known as quantum optics, which basically explores the nature and effects of quantized light. In this context, the most elementary quantum model of light–matter interaction consists of a highly idealized atom (modeled as a two-level system) coupled to a single mode of the quantized EM field. Under the rotating wave approximation (RWA), this model, known as the Jaynes–Cummings (JC) model [1, 2], remarkably admits an exact analytical solution, and despite the simplifying assumptions underlying its construction [3], the JC model provides deep insights into several fundamental aspects of light–matter interaction. Originally developed to compare semiclassical and quantum approaches to spontaneous emission [4], the JC model has since evolved to uncover nonclassical phenomena such as experimentally observed collapses [5] and revivals of the Rabi oscillations (RO) [6, 7, 8, 9, 10], squeezing [11, 12], sub-Poissonian statistics [13, 14], atom-field entanglement [15, 16, 17], and the generation of Schrödinger cat states [18, 19]. Furthermore, the core of the model is a two-level system (TLS) coupled to a harmonic oscillator, which allows it to be mapped onto a trapped ion [20, 21, 22] and relativistic systems [23]. Regarding applications, the JC framework is used in several quantum information processing schemes [24, 25]. There are also generalizations of the model involving additional field modes and atomic levels [26, 27], multiple two-level atoms [28, 29], quantum deformations [30, 31, 32], Kerr-like media [33, 34] and external pumpings [35, 36] Besides, the model naturally extends to the description of open quantum systems, where interaction with the environment leads to dissipation and decoherence [37, 38, 39].
In 1996, the collapses and revivals of the RO were observed in a remarkable cavity quantum electrodynamics experiment by Haroche’s group [10]. Using Rydberg atoms as a TLS [40, 41], this achievement provided a direct experimental verification of the quantum nature of the EM field and formed part of the body of work that was later recognized with the 2012 Nobel Prize in Physics awarded to Serge Haroche [42] jointly with David Wineland because of his work on trapped ion systems [43, 44]. The JC model’s pivotal role in this achievement highlights its fundamental importance, a fact underscored by the exponential growth of interest in the system since then [2].
@@
A particularly interesting generalization is obtained when the model parameters are allowed to vary in time, giving rise to the time-dependent JC (TDJC) model [45, 24]. This may account for atomic motion [46, 47], transient effects in the cavity [48, 49], or the incorporation of stochastic aspects [50]. However, unlike the standard JC model, analytical methods are not always applicable. Notable exceptions are the Nikitin [48], Rosen-Zener [49], Landau-Zener [51, 52] and Demkov–Kunike [51] models. Remarkably, under resonance conditions, it is possible to obtain exact solutions considering a time-dependent atom-field coupling [46, 53, 47]. From an experimental point of view, Ref. [54] suggests that a time-dependent coupling may be achievable in cavity quantum electrodynamics through variations in the atom’s position or by changing an external electric field.
@@
The objective of this work is twofold: we first present a didactic overview of the solution of the resonant TDJC model in terms of the evolution operator. We discuss two specific cases of time-dependent atom–field coupling: (i) linear [53] and (ii) hyperbolic secant [49]. The latter is particularly desirable in quantum control [55], where manipulating the TLS is one potential objective [56, 57]. We focus on the dynamics of the atomic population inversion and the von Neumann entropy, examining the impact of the time dependence on both quantities. The subsystems are assumed to be initially prepared in pure states, in which case the entropy directly reflects the atom–field entanglement. In the second part of the paper, we consider the atom moving inside the cavity, which leads to a sinusoidal time dependence for the atom–field coupling. Related scenarios have been explored in the literature, particularly in studies of entanglement [58, 59], self-induced transparency effects [46], and the role of atomic coherence in the JC model with motion [60]. We also assume a different set of initial conditions, namely, the field in a thermal state and the atom prepared either in the excited state or in a superposition of atomic ground and excited states (eigenstate of $\hat{\sigma}_{x}$). The JC model with a partially cooled cavity is relevant, for example, for the study of entanglement [61, 59], and is also experimentally accessible [9]. We analyze the atomic evolution from a different perspective, focusing on the dynamics of the Bloch vector, and compare our results to the established findings [62, 61, 63, 64] within the context of constant coupling. We find that the behavior of the quantities studied, basically the components and modulus of the Bloch vector, is generally dominated by periodic coupling, even in the presence of thermal noise, in contrast to the case of constant coupling [61]. We note that a periodic evolution of both the field entropy and the atomic inversion, induced by atomic motion and in the absence of thermal noise, was reported in [47]. Other effects may arise, manifesting in the Bloch vector dynamics. Two control parameters in the system, $p$, the number of half-wavelengths of the standing-wave cavity, and $\zeta_{3}$, proportional to the velocity of the moving atom, have similar effects. For instance, smaller values of $p$ and $\zeta_{3}$ lead to a dynamical evolution with longer periods, more closely resembling the behavior of the constant-coupling case. Interestingly, increasing the thermal fluctuations, which corresponds to a larger mean photon number in the initial field state, leads to a form of atomic population trapping [62], while the periodicity of the Bloch vector dynamics tends to persist despite the added noise.
@@ III Time-dependent coupling dynamics
In this work, we are interested in how time-dependent coupling parameters affect the on-resonance dynamics of the JC model. Other works that adopt a similar approach can be found in Refs. [46, 95, 53, 47, 96]. We remark that stochastic effects can be incorporated using a time-dependent coupling, as demonstrated in Ref. [50]. Therein, phase fluctuations are modeled via a random telegraph process to study the effects of decoherence. In this approach, the oscillations of both the atomic population inversion and the second-order correlation function of the field exhibit damping that is inversely proportional to the mean time interval between stochastic phase changes. While this is an interesting scenario, here we focus on well-behaved time-dependent couplings, and a detailed treatment of stochastic couplings is not addressed within the present analysis.
@@ III.1 Linear
Joshi and Lawande [53] considered the linear coupling modulation given by
@@
where $\lambda_{0}$ and $\zeta_{1}$ are control parameters that determine whether the coupling changes suddenly or adiabatically. According to [24], this modulation can be used to represent the scenario of a well-localized atom in a spatially varying mode. The authors observed that, for an initial coherent field state, this time-dependent coupling alters the revival time. More recently, Hu and Tan [97] studied a double JC model with linear modulation, focusing on the disappearance of entanglement between the two atoms. They suggested that a linear coupling parameter could be realized in cavity quantum electrodynamics by controlling the intensity of an external electric field.
@@ III.2 Hyperbolic Secant
This type of coupling can be used to simulate transient effects in the cavity, allowing the interaction to be switched on and off, depending on the choice of the initial time. The population inversion in this scenario is given by
@@
In Fig. 5, we observe how this coupling, which decays to zero, can lead to different asymptotic values for both population inversion and entanglement. In Fig. 5(a), we set $\zeta_{2}=0.3$ (faster decay of the coupling), causing the dynamics to stop before reaching the revival time. In contrast, in Fig. 5(b), we adopt a slower decay for the coupling. In both cases, the dynamics cease after the collapse, which causes the asymptotic value of $W(t)$ to be 0. However, for another combination of parameters, it would be possible to obtain different asymptotic values. Furthermore, another important aspect is that, even though the asymptotic value of the population inversion is the same in both cases, the von Neumann entropy value is significantly different. The coherence terms of the density matrix, which do not affect the atomic probabilities are influenced by the different atom-field decoupling times, and this is reflected in the entanglement behavior.
@@ IV Atomic motion
As mentioned in the previous section, Schlicher was the first to introduce a time-dependent coupling parameter. In that context, he used a sinusoidal modulation to account for the motion of the atom along the standing-wave [46]. Since then, this type of time dependence has been extensively studied: Bartzis investigated the squeezing in the cavity’s degree of freedom [95], Wilkens and Meystre extended the approach to maser systems [98], and Fang computed the von Neumann entropy and the Husimi function [47], considering an initial coherent state of the cavity. Other works onatomic motion in the model can be found in Refs. [99, 100, 101, 102, 103]. In this case, we adopt a semiclassical approach, in which atomic motion is treated classically – an approximation justified when the kinetic energy of the atom significantly exceeds the interaction energy of the JC model [24].
In cavity quantum electrodynamics experiments, the atom crossing the cavity goes through the spatial profile of the EM field mode, being subjected to a time-dependent coupling. To simplify the analysis of the experiments, it is introduced a constant effective interaction time $t_{eff}$ obtained by averaging the spatial variation of the atom–field coupling as the atom crosses the cavity mode. Thus, the atom is treated as if it traversed the cavity in a time $t_{eff}$ inversely proportional to the atom’s velocity while experiencing the maximum atom–field coupling [104, 105, 78]. This is justified because, at resonance, the atom–field Hamiltonians at different times commute, and the time-ordered exponential of the evolution operator reduces to a simple exponential [78]. The precise control of the atom-field interaction time was essential for the first experimental observation of collapses and revivals, accomplished by Haroche’s group [10]. For a didactic overview of the experimental methods related to the JC model in electrodynamic cavities, we recommend Ref. [78]. In the present work, we treat the time-dependent coupling explicitly in the dynamical evolution of the system, as we are interested in the effects of such time-varying interaction.
@@ IV.1 Bloch vector
Figure 8: Parameter plot of the $R_{x}(t)$ and $R_{z}(t)$ components as a function of time. In panel (a), atomic motion is not considered, and we set the parameter $\lambda=1$. Conversely, in the remaining panels, atomic motion is included, with parameters $\lambda_{0}=1$, $v=\zeta_{3}\pi/L$ and considered a time interval $0<t<40$. In panels (b) and (c), we set $p=1$, whereas in panel (d), we use $p=2$. In (b), we employ $\zeta_{3}=0.25$, while in the other panels, we fix $\zeta_{3}=1$. In all cases, the initial average photon number is $\langle n\rangle=0.5$. We observe that atomic motion dominates the disordered behavior of the Bloch vector.
Figure 9: Time-evolution of the components of the Bloch vector: $R_{x}(t)$ (solid green line), $R_{z}(t)$ (dotted purple line), and $R(t)$ (dashed black line). In panel (a), atomic motion is not considered, and we set the parameter $\lambda=1$. Conversely, in the remaining panels, atomic motion is included, with parameters $\lambda_{0}=1$, $v=\zeta_{3}\pi/L$ and considered a time interval $0<t<40$. In panels (b) and (c), we set $p=1$, whereas in panel (d), we use $p=2$. In (b), we employ $\zeta_{3}=0.25$, while in the other panels, we fix $\zeta_{3}=1$. In all cases, the initial average photon number is $\langle n\rangle=0.5$. Additionally, an increase in $p$ diminishes the amplitude of $R_{x}(t)$ and $R(t)$.
@@
In Fig. 8, we plot the components $R_{x}(t)$ and $R_{z}(t)$ as functions of time in a parameter plot. In the constant coupling case, shown in panel (a) of Fig. 8, we set $\lambda=1$. For the sinusoidal coupling case, depicted in the other panels of Fig. 8, we fix the parameters as $\lambda_{0}=1$ and $v=\zeta_{3}L/\pi$. In panels (b) and (c) we set $p=1$, while in panel (d), we set $p=2$. In (b), we employ $\zeta_{3}=0.25$, while in the other panels, we fix $\zeta_{3}=1$. At the outset, we choose a low average photon number, $\langle n\rangle=0.5$, corresponding to a low temperature, for all scenarios. Comparing panels (a) with (b), (c) and (d) of Fig. 8, we observe that the sinusoidal coupling introduces periodicity into the Bloch vector’s motion, replacing the disordered behavior. In contrast, for a smaller $\zeta_{3}$, as shown in (b), the behavior partially resembles that of the time-independent coupling case, exhibiting reduced-amplitude oscillations within each period. This suggests that for lower values of the parameter $\zeta_{3}$, an intermediate regime emerges in which vestiges of the aperiodic evolution persist.
In Fig. 9, we present the time-evolution of the components and absolute value of the Bloch vector for both constant and sinusoidal coupling scenarios. The parameters in Fig. 9(a)-(d) correspond to those in Fig. 8(a)-(d), respectively. In Fig. 9(b), (c) and (d), we again observe the periodicity inherited from the modulated coupling, as compared to Fig. 9(a). Furthermore, as shown in panels (b) and (c), the increase to $\zeta_{3}=1$ results in a higher frequency of the oscillations. Additionally, the amplitudes of the oscillations in $R_{x}(t)$ decrease, whereas $R_{z}(t)$ oscillates with approximately the same amplitude.
Furthermore, a comparison of Figs. 9(c) and (d) shows that an increase in the number of half wavelengths, $p$, not only leads to faster oscillations, but also reduces the amplitude of $R_{x}(t)$ and $R(t)$, while notably increasing the time-averaged values of these quantities. This is because a larger $p$ favors the alignment of the atom with the field, thus increasing the mean value of $R_{x}(t)$ and also the mean purity of the atomic state, here quantified by $R(t)$. We also note a subtle decrease in the amplitude of the oscillations in $R_{z}(t)$, indicating a weaker influence of the cavity on the atomic populations. This effect can also be attributed to the enhanced alignment of the atom with the field.
In all examples with atomic motion, we begin with $R(0)=R_{x}(0)=1$ and $R_{z}(0)=0$, meaning that initially, the atomic subsystem is pure, the atomic transition is aligned with the field, and the excited and ground populations are equal. The evolution of the system leads to a periodic return to the initial state. This behavior is a direct consequence of the trigonometric dependence of the coupling area $A(t)$, as discussed previously, which can also be interpreted in terms of Bloch vector rotations for the chosen velocity, as predicted by Joshi and Xiao [60].